A student calculated the molarity of a solution prepared by dissolving 0.730 mol of table sugar (sucrose, C12H22O11) in 1.8x10^3 mL of water as 4.06x10^-4 M C12H22O11


1
Expert's answer
2021-05-26T07:04:30-0400

"V=1.8 \times 10^3 \;mL = 1.8 \;L \\

n = 0.730 \;mol"

Proportion:

0.730 mol – 1.8 L

x mol – 1.0 L

"x = \frac{0.730 \times 1.0}{1.8}=0.4055\;M"

Student’s answer is wrong.

The right answer is 0.4055 M.